Bank has a net profit margin of 5.6%, an asset utilization ratio of 18%, and an equity multiplier of 15. Calculate the bank's ROA & ROE.
Answer #1

Net Profit Margin = Net Income/Sales = 0.056

Asset Utilization ratio = Sales / Total Assets = 0.18

Equity Multiplier = Total Assets / Total Equity = 15

ROA = Net Income / Total Assets = Asset Utilization Ratio * Net Profit Margin = 0.18*0.056 = 0.01008

ROE = Net Income / Total Equity = Asset Utlization Ratio * Equity Multiplier * Net Profit Margin = 0.18*15*0.056 = 0.1512